# Binary Decisions in DAOs: Accountability and Belief Aggregation via Linear Opinion Pools

### 10 Apr 2026 v2

@@ 1 Introduction
We identify two primary pathways for evolving DAO governance beyond its current limitations. The first, and the one pursued in this work, is the design of decision-making mechanisms for governance councils that ensure their members are directly accountable for the impact of their decisions on the broader community. The second focuses on developing community-wide decision-making mechanisms that enable effective information aggregation to reach better decisions. Notable examples of the latter include decision markets, as implemented by MetaDAO [metadao], and holographic consensus, as employed by DAOStack [daostack_whitepaper].
Kiayias et al. defines two essential properties regarding blockchain governance [kiayias2022sok]. Although their work focuses on blockchain governance—a subcategory of DAO —these properties are readily extendable to the broader DAO context.
@@
A DAO satisfies the property of Accountability if, whenever participants bring in a change, they are held individually responsible for it in a clearly defined way by the platform.
Definition 1.2 (Sustainable Participation (informal)).
A DAO sustains participation if it incentivizes, via monetary rewards or otherwise, participants who participate in the decision-making process of the platform.
These properties are in regard to the whole governance protocol, not the council’s internal decision-making mechanism. For example, according to the evaluation performed in that study, only Polkadot fully satisfies Accountability at the protocol level, as it does so by forcing users who vote in favor of a proposal to lock tokens until the proposal is deployed. A proposal is deployed if it is approved by the council. An unsolved issue, however, lies in the council itself: even if its members receive some direct compensation, they are not directly accountable for the changes they bring. This is because Polkadot uses majority voting as a decision-making mechanism in the council of experts. In Appendix A, we show that this mechanism does not satisfy Accountability.
These properties are presented as an overview in the original article, but they remain largely informal. Furthermore, they do not consider the decision-making mechanism used by the council. This paper bridges this gap by providing a more formal treatment of these properties regarding the decision-making mechanism in the councils. Sustainable participation is an informal statement of a known property in mechanism design theory: strict Individual Rationality. Accountability, by contrast, is seldom addressed within this area.
In an ideal decision-making mechanism, experts whose votes are pivotal—that is, whose votes actually change the outcome—should be rewarded when that outcome is positive and penalized when it is negative. To achieve this, mechanisms such as Decision Markets condition transfers on the observed outcome. This has been done in practice in DAOStack and MetaDAO [daostack_whitepaper, metadao].
@@ 1.1 Model Overview and Contributions
A key distinction arises from the interplay between these two types of private information. An expert is aligned when their preferences and beliefs point in the same direction, and unaligned when they conflict. This distinction is central to our results, as the mechanism’s incentive guarantees differ across the two cases. The full model is presented in Section 3.
Our contributions are:
@@
    We formalize Accountability and Sustainable Participation as mechanism design properties for governance councils, bridging the gap between the informal governance requirements identified in [kiayias2022sok] and classical mechanism design theory. We formalize the designer’s objective as a classification problem: align the decision with the aggregate expert beliefs rather than maximize the sum of idiosyncratic utilities.
@@ 2.4 Decision Making Mechanisms
Table 1 compares decision-making mechanisms along the following dimensions: whether the mechanism elicits experts’ subjective beliefs about the success of each alternative (Belief Elicitation); whether it mitigates the influence of experts’ private preferences on the collective decision (Reduce Idiosyncratic Noise); whether it supports multiple simultaneous participants (Multi-Agent); whether it operates in a single round (One-Shot); whether it functions without external funding (No Subsidies); whether total transfers sum to zero (Budget Balanced); and whether it operates without observing the realized outcome (No Outcome Observation).
@@ 3.5 Formalization of the Properties
Informally, this condition ensures that no expert is made worse off by participating in the mechanism. Their expected utility from engaging and submitting their optimal message $m_{i}^{\star}$ must be at least as great as their expected utility from abstaining, given their private type $\tau_{i}$ and the expected strategies of all other experts $\mathbf{m}_{-i}$.
Furthermore, a mechanism satisfies Strict Interim Individual Rationality if this inequality is strict. This strict version of Interim IR serves as the formal definition for Sustainable Participation. It guarantees that experts derive a strictly positive expected net benefit from participating in the decision-making mechanism. Note that not participating in the mechanism does not yield 0 utility to the expert since they will be affected by the outcomes of the decision. This includes the reward rules, that are transferred independently of the participation of the expert in the mechanism, i.e., the transfers are applied just because the expert is a member of the council.
@@ 4.1 Characterizing the Reward Rule
Before constructing the mechanism, we give an observation that serves as its building block. {observation}[Additive Adjustments] Consider a VCG mechanism as in Section $3.$ Suppose that, in addition to the VCG transfer, each expert $i$ receives a fixed reward $r\in\mathbb{R}$ whenever a specific alternative is selected, where $r$ does not depend on the submitted messages. Then a risk-neutral expert’s dominant strategy shifts from reporting their true valuation true preference intensity $v_{i}$ to:
@@ 7     $a\leftarrow B$;
10
11    Observe the realized outcome $\Delta\in\{-1,1\}$;
12
@@ 21          $\pi_{i}\leftarrow t_{i}+r_{i}$;
24          return Allocation $a$, payoffs $\boldsymbol{\pi}=(\pi_{1},\ldots,\pi_{n})$;
@@ 5.2 Incentive Properties
Under mechanism $\mathcal{M}$, for any expert $i$ (aligned or unaligned), every deviation from $m_{i}^{\star}$ that changes the allocation toward the alternative with lower success probability from the expert’s own perspective is weakly dominated. Formally, any $m_{i}^{\prime}\neq m_{i}^{\star}$ such that the allocation changes from $b$ to $a$ with $p_{i}^{a}<p_{i}^{b}$ satisfies:
@@
Case expert pivots toward their preference (Outcome is AA). Here $a=A$ and $b=B$ and $m_{i}^{\star}\geq 0$ and $t_{i}=\sum_{j\neq i}m_{j}\geq-m_{i}^{\star}$. The IR condition reduces to $t_{i}\geq-m_{i}^{\star}$, which is exactly the pivotality constraint, and thus it must hold.
@@
The inequality is strict whenever $t_{i}>-|m_{i}^{\star}|$. If experts are uncertain about the reports of others and assign positive probability to being pivotal, the inequality holds strictly in expectation. Sustainable Participation follows: experts derive a strictly positive expected net benefit from participating.
@@
The belief signal can be written as $B(\tau)=\sum_{i=1}^{n}w_{i}(p_{i}^{A}-p_{i}^{B})$ with strictly positive weights. Since $x^{*}(\mathbf{p})$ selects AA if and only if $\sum w_{i}(p_{i}^{A}-p_{i}^{B})>0$, we have $\mathrm{sgn}(B(\tau))=\mathrm{sgn}(x^{*}(\mathbf{p}))$.
@@
The mechanism performs best when experts’ beliefs converge. Each term in $N(\tau)$ carries the sign of the expert’s preference direction, but the denominators $2(1-p_{i}^{(\cdot)})$ depends on belief strength. As experts’ beliefs converge toward the same alternative — say all experts come to share the belief that $p_{i}^{B}>p_{i}^{A}$ — experts who prefer AA see their weight shrink, which reduces their positive contribution to $N(\tau)$. At the same time, experts who prefer $B$ see their weight grow, amplifying their negative contribution. This leads $N(\tau)$ to be driven towards the same sign as $B(\tau)$. When $\mathrm{sgn}(N)=\mathrm{sgn}(B)$, the threshold $\bar{c}(\tau)=0$ by (9), and the mechanism classifies correctly for any budget $c\geq 0$. This convergence property makes the mechanism particularly suited to DAO governance, where proposals are typically preceded by community discussions and temperature checks that foster belief alignment among experts before the formal vote.
Figure 2 illustrates the classification plane. Every type profile $\tau$, regardless of the number of experts $n$, maps to a single point $(N(\tau),\,B(\tau))$. The decision boundary $N+\tfrac{c}{n}B=0$ is a line through the origin with slope n/c-n/c; increasing the per-expert budget $\tfrac{c}{n}$ rotates this line toward the NN-axis, shrinking the misclassification wedge (shaded red for $c_{2}/n$). In the green quadrants noise and signal agree in sign, so $\bar{c}(\tau)=0$ and the mechanism classifies correctly for any budget. A point’s distance from the NN-axis relative to its distance from the $B$-axis determines its threshold $\bar{c}(\tau)=|N|/|B|$: points closer to the NN axis (strong noise, weak signal) are harder to classify and require a larger budget.
@@ 5.4 Example
We illustrate the mechanism with a concrete scenario. A DAO governance council of $n=3$ experts must decide whether to integrate a new DeFi protocol (AA = approve) or reject it ($B$ = reject). The DAO allocates a budget of $c=24$, yielding $c/n=8$ per expert. Table 2 summarizes the experts’ private types.
@@
Expert 1 2 0 0.75 0.50 Aligned (favors AA)
@@
Expert 3 1 0 0.50 0.75 Unaligned (prefers AA, believes $B$)
@@
Expert 3 submits a negative report ($m_{3}^{\star}=-3$) despite personally preferring AA. The belief component (4-4) outweighs the preference component (+1+1), pulling the report toward $B$ — the alternative Expert 3 believes will benefit the organization.
@@
The noise $N(\tau)=3>0$ pushes toward AA (driven by idiosyncratic preferences), while the belief signal $B(\tau)=-0.5<0$ pushes toward $B$. Since $\mathrm{sgn}(B)<0$, the designer’s optimal decision is $B$. The aggregate confirms: $N+\tfrac{c}{n}B=3+8(-0.5)=-1<0$, correctly selecting $B$. The budget threshold is $\bar{c}(\tau)=|N|/|B|=6$, and indeed $c/n=8>6$.
@@
With a reduced budget $c=12$ ($c/n=4<6=\bar{c}$), the reports become $m_{1}^{\star}=8$, $m_{2}^{\star}=-6$, $m_{3}^{\star}=-1$, yielding $\sum m_{i}^{\star}=1>0$. The mechanism selects AA — a misclassification, because the budget is insufficient to overcome the preference-driven noise.
@@
Suppose Expert 3 deviates from $m_{3}^{\star}=-3$ to $m_{3}^{\prime}=5$, following their personal preference for AA. Since $\sum_{j\neq 3}m_{j}=2>0$ and $m_{3}^{\prime}+2=7>0$, Expert 3 is no longer pivotal ($t_{3}=0$) and the mechanism selects AA. Expert 3’s expected utility drops from 33 (at $m_{3}^{\star}$) to 11 (at $m_{3}^{\prime}=5$): the deviation toward the alternative Expert 3 believes is less likely to succeed is costly, confirming Proposition 5.8.
@@ 6 Conclusions
We presented a decision-making mechanism for governance councils in DAOs that augments VCG transfers with outcome-contingent rewards. Unlike classical mechanism design, where the objective is to maximize social welfare, we frame decision-making as a classification problem: the designer seeks to align the collective decision with the aggregate expert beliefs. The mechanism satisfies Budget Constraint, Symmetry, Belief Neutrality, Weak Accountability, and Interim Individual Rationality. For aligned experts it is DSIC; for unaligned experts, the Safe Deviation property guarantees that no expert can profitably deviate from $m_{i}^{\star}$ toward an alternative they individually believe is less likely to succeed. The main result (Theorem 5.12) decomposes the aggregate signal into idiosyncratic noise and a belief signal whose sign matches the designer’s optimal decision, with correct classification achieved whenever the per-expert budget exceeds a type-dependent threshold $\bar{c}(\tau)$. This threshold decreases as experts’ beliefs converge — a condition naturally fostered by deliberative processes that precede governance votes in DAOs.

### 13 Apr 2026 v3

@@ 1 Introduction
Since governance councils are already the predominant decision-making structure in large DAOs, the question is how the council itself should decide. The council’s decision-making mechanism should elicit and aggregate the experts’ beliefs about which alternative best serves the organization, rather than let their idiosyncratic preferences determine the outcome. To achieve this, experts need skin in the game, their payoffs must depend on the realized outcome of the decision, so that an expert who bears the consequences of a bad decision reports according to their genuine beliefs. Decision markets, as implemented by MetaDAO [metadao] and DAOStack [daostack_whitepaper], already condition transfers on the observed outcome, but they are designed for community-wide participation, not for councils. Kiayias and Lazos [kiayias2022sok] identify this requirement as Accountability:
@@
A DAO satisfies the property of Accountability if, whenever participants bring in a change, they are held individually responsible for it in a clearly defined way by the platform [kiayias2022sok].
Although the original work focuses on blockchain governance—a subcategory of DAO governance—this property is readily applicable to the broader DAO context. This property is in regard to the whole governance protocol, not the council’s internal decision-making mechanism. For example, according to the evaluation performed in that study, only Polkadot fully satisfies Accountability at the protocol level, as it does so by forcing users who vote in favor of a proposal to lock tokens until the proposal is deployed. A proposal is deployed if it is approved by the council. An unsolved issue, however, lies in the council itself: even if its members receive some direct compensation, they are not directly accountable for the changes they bring. This is because Polkadot uses majority voting as a decision-making mechanism in the council of experts. In Appendix A, we show that this mechanism does not satisfy Accountability.
Accountability is presented informally in the original article and is evaluated at the level of the whole governance protocol, not the council’s internal decision-making mechanism. Since accountability is necessary to align the experts’ reports with the organization’s interests, this paper formalizes it for the council setting. Our formal condition, Weak Accountability (Section 3.5), requires that an expert whose report changes the outcome receives a strictly lower total monetary transfer than a non-pivotal expert when the outcome is bad, and no lower when the outcome is good.
@@ 1.1 Model Overview and Contributions
A key distinction arises from the interplay between these two types of private information. An expert is aligned when their preferences and beliefs point in the same direction, and unaligned when they conflict. This distinction is central to our results, as the mechanism’s incentive guarantees differ across the two cases. The full model is presented in Section 3. Our contributions are:
@@
    Motivated by the informal property of Accountability identified in [kiayias2022sok], we define Weak Accountability (Section 3.5) for the council setting, and show that none of the standard decision-making mechanisms satisfies it (Appendix A).
@@ 2.2 Decision Markets and Scoring Rules
2.3 Two-Stage and Outcome-Contingent Mechanisms
Standard mechanism design determines transfers at the time of the decision, before the outcome is known. A separate line of work allows transfers to depend on information that becomes available after the decision is implemented, expanding the set of achievable allocations and incentive properties.
In auctions, [hansen1985auctions] showed that conditioning the payment on the realized value of the good (e.g., a share of future revenue) allows the seller to raise more revenue than under standard sealed-bid formats.
[mezzetti2004mechanism] generalized this idea to multi-agent mechanism design with interdependent valuations, where standard Groves mechanisms fail to achieve efficiency. Mezzetti proposed a two-stage Groves mechanism: in the first stage, agents report their types and the efficient allocation is selected; after the outcome is realized and agents observe their own payoffs, a second reporting stage collects these payoffs, and final transfers are computed from both stages. By making each agent the residual claimant of the total surplus, truthful reporting is a best response in both stages.
Our mechanism shares the two-stage architecture: VCG transfers are computed at the time of the decision, and an outcome-contingent reward is distributed after the evaluation period. It departs from this literature in two ways. First, the objective is not efficiency but classification: we seek to align the decision with the aggregate expert beliefs, not to maximize the sum of utilities. Accordingly, the outcome-contingent reward is designed to satisfy Weak Accountability, not to make agents residual claimants of the surplus. Second, the second stage does not require agents to report their realized payoffs, instead, a single boolean $\Delta\in\{-1,1\}$ is observed.
@@ 2.5 Decision Making Mechanisms
Table 1 compares decision-making mechanisms along the following dimensions: whether the mechanism elicits experts’ subjective beliefs about the success of each alternative (Belief Elicitation); whether it mitigates the influence of experts’ private preferences on the collective decision (Reduce Idiosyncratic Noise); whether it supports multiple simultaneous participants (Multi-Agent) and whether it operates in a single round (One-Shot).
@@ 3.5 Formalization of the Properties
Note that not participating in the mechanism does not yield 0 utility to the expert since they will be affected by the outcomes of the decision. This includes the reward rules, that are transferred independently of the participation of the expert in the mechanism, i.e., the transfers are applied just because the expert is a member of the council.
@@ 4.1 Characterizing the Reward Rule
Before constructing the mechanism, we give an observation that serves as its building block. {observation}[Additive Adjustments] Consider a VCG mechanism as in Section 3.3. Suppose that, in addition to the VCG transfer, each expert $i$ receives a fixed reward $r\in\mathbb{R}$ whenever a specific alternative is selected, where $r$ does not depend on the submitted messages. Then a risk-neutral expert’s dominant strategy shifts from reporting their true valuation true preference intensity $v_{i}$ to:
@@ 7     $a\leftarrow B$;
10    for $i=1,\dots,n$ do
11       $t_{i}\leftarrow\min\!\left\{\,0,\;\,-\,\!\bigg|\,\sum_{j\neq i}m_{j}\,\bigg|\,\right\}$
12       end for
13      return Allocation $a$ and transfers $\mathbf{t}=(t_{1},\ldots,t_{n})$;
14
15       Stage 2: Ex-post reward Observe the realized outcome $\Delta\in\{-1,1\}$;
@@ 21             $r_{i}\leftarrow t_{i}-\frac{c}{n}$;
26            return Rewards $\mathbf{r}=(r_{1},\ldots,r_{n})$;
@@ 5.2 Incentive Properties
Under mechanism $\mathcal{M}$, for any expert $i$, every deviation from $m_{i}^{\star}$ that changes the allocation toward the alternative with lower success probability from the expert’s own perspective is weakly dominated. Formally, any $m_{i}^{\prime}\neq m_{i}^{\star}$ such that the allocation changes from $b$ to $a$ with $p_{i}^{a}<p_{i}^{b}$ satisfies:
@@
Case expert pivots toward their preference (Outcome is $A$). Here $a=A$ and $b=B$ and $m_{i}^{\star}\geq 0$ and $t_{i}=\sum_{j\neq i}m_{j}\geq-m_{i}^{\star}$. The IR condition reduces to $t_{i}\geq-m_{i}^{\star}$, which is exactly the pivotality constraint, and thus it must hold.
@@
The belief signal can be written as $B(\tau)=\sum_{i=1}^{n}w_{i}(p_{i}^{A}-p_{i}^{B})$ with strictly positive weights. Since $x^{*}(\mathbf{p})$ selects $A$ if and only if $\sum w_{i}(p_{i}^{A}-p_{i}^{B})>0$, we have $\mathrm{sgn}(B(\tau))=\mathrm{sgn}(x^{*}(\mathbf{p}))$.
@@
The mechanism performs best when experts’ beliefs converge. Each term in $N(\tau)$ carries the sign of the expert’s preference direction, but the denominators $2(1-p_{i}^{(\cdot)})$ depends on belief strength. As experts’ beliefs converge toward the same alternative — say all experts come to share the belief that $p_{i}^{B}>p_{i}^{A}$ — experts who prefer $A$ see their weight shrink, which reduces their positive contribution to $N(\tau)$. At the same time, experts who prefer $B$ see their weight grow, amplifying their negative contribution. This leads $N(\tau)$ to be driven towards the same sign as $B(\tau)$. When $\mathrm{sgn}(N)=\mathrm{sgn}(B)$, the threshold $\bar{c}(\tau)=0$ by (9), and the mechanism classifies correctly for any budget $c\geq 0$. This convergence property makes the mechanism particularly suited to DAO governance, where proposals are typically preceded by community discussions and temperature checks that foster belief alignment among experts before the formal vote.
Figure 2 illustrates the classification plane. Every type profile $\tau$, regardless of the number of experts $n$, maps to a single point $(N(\tau),\,B(\tau))$. The decision boundary $N+\tfrac{c}{n}B=0$ is a line through the origin with slope n/c-n/c; increasing the per-expert budget $\tfrac{c}{n}$ rotates this line toward the $N$-axis, shrinking the misclassification wedge (shaded red for $c_{2}/n$). In the green quadrants noise and signal agree in sign, so $\bar{c}(\tau)=0$ and the mechanism classifies correctly for any budget. A point’s distance from the $N$-axis relative to its distance from the $B$-axis determines its threshold $\bar{c}(\tau)=|N|/|B|$: points closer to the $N$ axis (strong noise, weak signal) are harder to classify and require a larger budget.
@@ 5.4 Example
We illustrate the mechanism with a concrete scenario. A DAO governance council of $n=3$ experts must decide whether to integrate a new DeFi protocol ($A$ = approve) or reject it ($B$ = reject). The DAO allocates a budget of $c=24$, yielding $c/n=8$ per expert. Table 2 summarizes the experts’ private types.
@@
Expert 1 2 0 0.75 0.50 Aligned (favors $A$)
@@
Expert 3 1 0 0.50 0.75 Unaligned (prefers $A$, believes $B$)
@@
Expert 3 submits a negative report ($m_{3}^{\star}=-3$) despite personally preferring $A$. The belief component (4-4) outweighs the preference component (+1+1), pulling the report toward $B$ — the alternative Expert 3 believes will benefit the organization.
@@
The noise $N(\tau)=3>0$ pushes toward $A$ (driven by idiosyncratic preferences), while the belief signal $B(\tau)=-0.5<0$ pushes toward $B$. Since $\mathrm{sgn}(B)<0$, the designer’s optimal decision is $B$. The aggregate confirms: $N+\tfrac{c}{n}B=3+8(-0.5)=-1<0$, correctly selecting $B$. The budget threshold is $\bar{c}(\tau)=|N|/|B|=6$, and indeed $c/n=8>6$.
@@
With a reduced budget $c=12$ ($c/n=4<6=\bar{c}$), the reports become $m_{1}^{\star}=8$, $m_{2}^{\star}=-6$, $m_{3}^{\star}=-1$, yielding $\sum m_{i}^{\star}=1>0$. The mechanism selects $A$ — a misclassification, because the budget is insufficient to overcome the preference-driven noise.
@@
Suppose Expert 3 deviates from $m_{3}^{\star}=-3$ to $m_{3}^{\prime}=5$, following their personal preference for $A$. Since $\sum_{j\neq 3}m_{j}=2>0$ and $m_{3}^{\prime}+2=7>0$, Expert 3 is no longer pivotal ($t_{3}=0$) and the mechanism selects $A$. Expert 3’s expected utility drops from $3$ (at $m_{3}^{\star}$) to $1$ (at $m_{3}^{\prime}=5$): the deviation toward the alternative Expert 3 believes is less likely to succeed is costly, confirming Proposition 5.8.
@@ 6 Conclusions
We presented a decision-making mechanism for governance councils in DAOs that augments VCG transfers with outcome-contingent rewards. Unlike classical mechanism design, where the objective is to maximize social welfare, we frame decision-making as a classification problem: the designer seeks to align the collective decision with the aggregate expert beliefs. The mechanism satisfies Budget Constraint, Symmetry, Belief Neutrality, Weak Accountability, and Interim Individual Rationality. For aligned experts it is DSIC; for unaligned experts, the Safe Deviation property guarantees that no expert can profitably deviate from $m_{i}^{\star}$ toward an alternative they individually believe is less likely to succeed. The main result (Theorem 5.12) decomposes the aggregate signal into idiosyncratic noise and a belief signal whose sign matches the designer’s optimal decision, with correct classification achieved whenever the per-expert budget exceeds a type-dependent threshold. This threshold decreases as experts’ beliefs converge — a condition naturally fostered by deliberative processes that precede governance votes in DAOs.